OpenStudy (anonymous):

help me write this expression in factored form: x^2-4x+1

OpenStudy (anonymous):

it does not factor over the integers, but if you find the two zeros \(r_1\) and \(r_2\) then you can factor it as \[(x-r_1)(x-r_2)\]

OpenStudy (anonymous):

u can use quadratic formula x = (-b +- sqrt(b² -4ac)) / 2a or x^2 - 4x +1 = 0 (I'm assuming) x^2 -4x =-1 x^2-4x +4 = 3 (x-2)^2 = 3 x-2 = + or - square root(3) x= 2+ or - square root (3)